Evaluate the integral by making the given substitution. (Use C for the constant of integration.) *? Vx2 + 7 dx, u = x3 + 7
Answer #1

we are given

\int x^2\sqrt{x^3+7}dx

we can use u-subs

u=x^3+7

du=3x^2dx

=\int \frac{\sqrt{u}}{3}du

=\frac{1}{3}\cdot \int \:u^{\frac{1}{2}}du

we can use power rule

=\frac{1}{3}\cdot \frac{u^{\frac{1}{2}+1}}{\frac{1}{2}+1}

now, we can plug back u

and we get

=\frac{2}{9}\left(x^3+7\right)^{\frac{3}{2}}+C.............Answer